User Tools

Site Tools


ao:diary

Diary

6 Dec 2015
- Theo, Mike, John discussed big picture strategy (TODO list), then tried to follow the standard alignment.
- Discovered beam shear, then found large (2-3cm) pupil misalignment and decided to re-align LABAO from scratch. Alignment procedure followed was:

  1. Center beam on spherical mirror. This wasn't done because it was fine.
  2. Center beam on LDM. Note that the target is inadequate, as it can shift left-right etc!
  3. Point beam parallel to the table at BRT secondary. Note that this requires the LDM to be at the right height! The tolerance in parallelism to the table is relatively loose… about 0.1 degrees. (at least 0.05 degrees was achieved).
  4. Move the secondary to centre the beam on the rail target.
  5. Move the primary to align the beams down the rail. Note that this movement was extremely small

- The two-hole mask test was difficult because there was only one two-hole mask! This is essential for all telescopes for the current procedure.
- No laptop or power supply available for the beacon focusing camera. This needs to live in a standard place and be part of the manual.
- Beacon focus was +1,500,000 counts out (i.e. way out by eye) and when it was in focus it looked perfect on the LABAO WFS… except that only half the pupil was illuminated.
- Discovered dust on the fiber tip and also vignetting probably due to no POP realignment done after the pupil shift.

7 Dec 2015
- Theo and Judit again realigned the S2 labao from scratch. This time the beam was parallel to the optical table throughout.
- S1 and S2 beacons focused. Mike realised that the dichroic was installed backwards. The pupil shift seen (beacon with YSO dichroic versus M10 alignement) is consistent with this error.
- Mike and Judit aligned the lab WFS for S1 (see alignment page). Due to mounting convenience, the labao WFS is conjugated to 120m upstream rather than the 145m upstream for the LDM.

Plans for tomorrow
- Dichroic sine wave alignment testing (MJI)
- Go through list
- Reverse dichroics

9 Dec 2015
- The pupil shift seen earlier in the week continues. Specifically… with a good M10 alignment and standard labao alignment, the beacon pupil is shifted on the labao WFS. When correct on LABAO, there was about a 10% error on M10. There are 3 possible reasons for this:

  1. The beacon pupil at the telescope isn't aligned with the centre of the M5 mirror.
  2. The laser pupil in the lab isn't aligned with the WFS pupil. NB this can only be accurately checked with the Blue LED.
  3. The boresight laser definition on the beacon and laser do not match - this can also be accurately checked with the Blue LED and a lab corner cube. Note that an on-sky 1 arcsec angular uncertainty corresponds to approximately 1.3mm in the 19mm beam. About 7% or half a lenslet.

- ueyed doesn't have any special username etc permissions, i.e.:
[labao ROOT:807] id mireland
uid=557(mireland) gid=500(chara) groups=500(chara)
[labao ROOT:808] id ueyed
uid=495(ueyed) gid=501(ueye) groups=501(ueye)
[ So it remains totally weird that only mireland can run labao on S1 especially ]
- Nearly (not quite) 0.5 lenslets shift on the LABAO wavefront sensor (point 2 above). So this explains part of the issue.
- The autocollimation procedure for S2 also had a shift of “0.2” boxes when looking at the green LED off the corner cube. This could be an issue with the corner cube! (or a green versus blue issue). But in any case it is another uncertainty of the right magnitude.

ao/diary.txt · Last modified: 2018/07/07 13:05 by jones